OpenStudy (petlovers123):

Ralph needs 6 cups of milk to make a chocolate shake for his family. His measuring device measures liquids by the pint. If 1 pint is equal to 2 cups, how many pints should Ralph measure? 3 pints 4 pints 8 pints 12 pints

OpenStudy (petlovers123):

@AAAlive

OpenStudy (petlovers123):

will medal for first ANSWER

OpenStudy (petlovers123):

@NotSoGoodAtMath-_-

ganeshie8 (ganeshie8):

2 cups = 1 pint 6 cups = how many pints ?

OpenStudy (lolacole12):

8 i think

OpenStudy (lolacole12):

oops 12

OpenStudy (petlovers123):

12 pints

OpenStudy (lolacole12):

yes thats right cuase six times 2 is 12

ganeshie8 (ganeshie8):

think again

ganeshie8 (ganeshie8):

don't wry about math use a bit of common sense

OpenStudy (lolacole12):

wait its not

OpenStudy (petlovers123):

um ok

OpenStudy (lolacole12):

3

OpenStudy (petlovers123):

3

ganeshie8 (ganeshie8):

sure ?

OpenStudy (petlovers123):

yes i guess

OpenStudy (lolacole12):

right because f 2=1 and 2 x 3 is 6 so 2=1+2=1+2=1 so 3 yay

ganeshie8 (ganeshie8):

good, then 3 is correct !

OpenStudy (lolacole12):

yes im learning lol

OpenStudy (petlovers123):

im going to wait for @ganeshie8

OpenStudy (petlovers123):

yay got it

ganeshie8 (ganeshie8):

good job

OpenStudy (lolacole12):

Good job

OpenStudy (anonymous):

1= 2 x 3 = 6 = 3 pints sorry im so late

OpenStudy (petlovers123):

its ok but thnx anyway